Home‎ > ‎Crew 165‎ > ‎

0311-DailySummaryReport

Hi Mission Support!

Daily Summary Report (DSR) for 10 March 2016!
SOL: 5
Person filling out Report: Ken Sullivan
Summary Title: Green Hab & Toilet
Mission Status: All systems nominal

Sol 5 Activity Summary: Assembled several shelving units in the GreenHab and cleaned it up. Toilet assessment for small leak in flapper valve. Did get fuel for ATV's today.

Anomalies in work: None 
Weather: Overcast, light wind.  
Look ahead plan: EVA#4 for Heidi Beemer's Thesis  
Crew Physical Status: Good!

 EVA: No EVA today. 
 Reports to be filed: - Commander
 Report- Engineering Report- Journalist
 Report- Daily Summary Report-
 Astronomy- EVA #4 request
 Support Requested: Teflon tape.